# AudioMoG: Guiding Audio Generation with Mixture-of-Guidance

### v1

@@ 3.1 Analysis
Motivation. As discussed above, previous guidance methods, CFG and AG, guide the sampling process with diverse principles, both showing advantages over unguided generation. CFG enhances consistency with conditional information, and AG removes dispersion by mitigating errors. The effectiveness of CFG has been extensively validated in audio generation across various data representations (Liu et al., 2023; Evans et al., 2025), conditional modalities such as text (Deepanway et al., 2023; Huang et al., 2023a), video (Luo et al., 2023; Xu et al., 2024), and network architectures (Liu et al., 2023; Evans et al., 2025; Li et al., 2024a; Hung et al., 2024). However, CFG can still yield suboptimal results due to its overemphasis on condition information. Particularly, it may miss relevant sound events or fail to accurately generate uncommon audio events. Recent work ETTA (Lee et al., 2025) has explored AG on T2A generation, showing increased generation diversity but observing strong sensitivity to the choice of weak model. Given that either CFG or AG has shown quality improvement for audio generation, while both consider a single guiding principle, we explore a mixture-of-guidance framework, aiming at composing guidance methods to fulfill stronger results by exploiting their complementary advantages, e.g., cumulative benefits, even without sacrificing sampling efficiency.
3.2 Framework
General setting. AudioMoG presents a mixture-of-guidance strategy as follows, involving $M$ guidance methods:
$\bm{\epsilon}_{\text{MoG}}(\bm{z}_{t},t,\bm{c})=\sum_{i=1}^{N}w_{i}\bm{\epsilon}_{i}(\bm{z}_{t},t,\bm{c}),\quad\text{s.t.}\sum_{i=1}^{N}w_{i}=1,$ (8)
where $\bm{\epsilon}_{i}$ is a denoiser network and $w_{i}\in\mathbb{R}$ is the corresponding weight. When $M=1$, AudioMoG considers a single guidance method, which extrapolates two denoising results as mentioned in (Karras et al., 2024a). Given $M\geq 2$, AudioMoG starts exploiting the complementary advantages of different guidance methods. However, this inevitably increases the complexity, which may result in longer inference time, as the framework considers both the terms required by different methods and the additionally produced interaction terms, e.g., the unconditional and bad term when combining CFG and AG. Hence, even though a MoG framework may yield stronger performance by considering more guiding principles and leveraging their complementary benefits, an essential consideration is the balance between synthesis quality and inference speed.
To improve guided audio generation, we empirically observe that combining two distinctive guidance methods, both of which have proven more advantageous than unguided conditional sampling, shows potential to achieve improved synthesis quality without sacrificing inference speed.
Hierarchical guidance. When considering a mixture of two guidance methods, the MoG framework shown by Equation 8 exploits their complementary advantages by linearly combining four noise predictors (i.e., $M=2$, $N=4$), which can be viewed as hierarchically combining two guidance methods (denoted by HG). Taking CFG and AG as examples for guided audio generation, AudioMoG can be interpreted from two perspectives  [^1] HG in CFG-AG or AG-CFG orders yields equivalent guidance family, as proven in Appendix A.. Firstly, we can interpret it as refining the CFG method with the AG guiding principle as follows:
@@ 3.2 Framework
Namely, the good and bad terms under the AG framework (Karras et al., 2024a) have been first enhanced by explicitly emphasizing condition alignment, and then AG further strengthens the result towards faithfully reconstructing the target distribution. Alternatively, as proven in Appendix A, it can be interpreted as applying the CFG method to more accurate score estimation results. Namely, the conditional and unconditional terms under the CFG framework (Ho & Salimans, 2022) have been first improved by AG to achieve higher score accuracy. This naturally outperforms previous works that solely depend on CFG: applying CFG to scores not yet refined by AG.
Other forms: parallel guidance, CFG, and AG. By controlling the weighting strategy for Equation 8, AudioMoG can show other forms exhibiting a different mechanism when exploiting the complementary advantages. When the interaction terms in Equation 8 are removed (i.e., $M=2$, $N=3$), AudioMoG degrades to employing CFG and AG in parallel (denoted by PG), which can be written as:
$\bm{\epsilon}_{\text{PG}}(\bm{z}_{t},t,\bm{c})=\bm{\epsilon_{\theta}}(\bm{z}_{t},t)+w_{1}(\bm{\epsilon_{\theta}}(\bm{z}_{t},t,\bm{c})-\bm{\epsilon_{\theta}}(\bm{z}_{t},t))+w_{2}(\bm{\epsilon_{\theta}}(\bm{z}_{t},t,\bm{c})-\bm{\epsilon}_{\bm{\theta}_{\text{bad}}}(\bm{z}_{t},t,\bm{c})).$ (10)
While PG qualitatively demonstrated by (Karras et al., 2024a) offers a simple and effective integration of CFG and AG, it implicitly assumes their compatibility at each sampling step. However, as discussed above, these two methods guide the model in potentially interfering directions: emphasizing the condition can lead to mode collapse, while correcting from weak references may introduce semantic drift. Thus, it may not be able to fully exploit their complementary advantages in comparison with exploiting the cumulative benefits, as demonstrated by experiments in Section 4.
It is also worth noting that, when further removing the terms in Equation 8, AudioMoG can recover the guidance method considering a single guiding principle, e.g., CFG or AG, indicating that MoG can be seen as a unified framework to incorporate diverse guidance methods.
Table 1: Objective metrics for text-to-audio generation on AudioCaps test set. The best performance for each metric is highlighted in bold, while the second-best is marked with an underline.
@@
AudioGen (Kreuk et al., 2022) 3.13 2.09 / / /
AudioGen-Large (Kreuk et al., 2022) 1.82 1.69 / / /
Make-An-Audio (Huang et al., 2023b) 1.61 1.61 7.29 18.32 /
TANGO-AF&AC-FT-AC (Kong et al., 2024) 2.54 / 11.04 17.19 /
AudioLDM-Large-Full (Liu et al., 2023) 1.96 1.59 8.13 23.31 0.43
AudioLDM 2 (Liu et al., 2024a) 2.09 1.79 8.14 26.44 0.50
AudioLDM 2-Large (Liu et al., 2024a) 1.89 1.54 8.55 26.18 0.53
Stable Audio Open (Evans et al., 2025) / 2.14 / / 0.35
CFG-only, $w=7$ 1.76 1.44 13.46 20.94 0.54
MoG-PG, $w_{1}=4.6,w_{2}=0.2$ 1.54 1.47 13.47 18.50 0.53
MoG-HG, $w_{1}=4.0,w_{2}=3.3,w_{3}=1.2$ 1.38 1.44 13.58 18.87 0.54
Table 2: Subjective metrics for text-to-audio generation on AudioCaps samples. Rated for overall quality and text relevance, with higher scores indicating better performance.
@@
4 Experiments
4.1 T2A experiment setup
Datasets. We use AudioSet (Gemmeke et al., 2017), FSD50k (Fonseca et al., 2021) and Clotho v2 (Drossos et al., 2020). To maintain consistency throughout the dataset, each track in these databases was segmented into 10-second clips and resampled at 16 kHz. The details of these datasets are further introduced in the Appendix F.1. To compare with prior work, we evaluated our models on the widely used AudioCaps benchmark (Kim et al., 2019), which consists of about 1K 10-second audio clips.
Model configurations. We use FLAN-T5(Chung et al., 2024) as the text encoder for our base model, and we train a Variational Autoencoder (VAE) (Kingma et al., 2013) that compresses the original waveform into the latent representation. A more detailed description of the model configurations and compression networks is provided in the Appendix F.3. We trained the model for 1 M iterations with a batch size of 8 per GPU. We used the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of 5e-5 and a condition drop $p_{\text{uncond}}=0.1$ for CFG. The bad model was trained under the same configuration as the main model, but with only 0.1 M iterations. During inference, we use DPM++ 2M SDE (Lu et al., 2022).
Evaluation metrics. We conduct a comprehensive evaluation of our models using both objective and subjective evaluations to assess audio generation quality, text-audio alignment, and inference efficiency. Objective metrics include Fréchet Audio Distance (FAD) (Kilgour et al., 2019), Kullback-Leibler (KL) divergence (Kullback & Leibler, 1951), Inception Score (IS) (Salimans et al., 2016), Fréchet Distance (FD) (Heusel et al., 2017), and LAION-CLAP score (Wu et al., 2023). For subjective evaluation, we recruited 20 human raters to score two aspects: (i) overall perceptual quality (OVL), and (ii) semantic relevance to the input text (REL). Both scores are rated on a 1–5 scale. More details are introduced in Appendix F.5 and F.6, respectively.
Table 3: Objective metrics for video-to-audio generation on VGGSound test set.
Model FAD $\downarrow$ KL $\downarrow$ IS $\uparrow$ FD $\downarrow$ IBS $\uparrow$ Align acc $\uparrow$
@@ 4.1 T2A experiment setup
IM2WAV (Sheffer & Adi, 2023) 6.41 2.54 / / 19.0 74.3
Diff-Foley (Luo et al., 2023) 5.79 3.12 10.8 21.90 20.4 89.9
FoleyGen (Mei et al., 2024) 1.65 2.35 / / 26.1 73.8
VTA-LDM (Xu et al., 2024) 2.01 2.37 10.4 12.80 26.2 77.0
FoleyCrafter (Zhang et al., 2024) 2.32 2.54 9.9 18.10 27.7 83.6
V2A-Mapper (Wang et al., 2024a) 0.90 2.68 12.5 8.35 22.4 78.3
VAB-Encodec (Su et al., 2024) 2.69 2.58 / / / /
VATT w/o text (Liu et al., 2024b) 2.35 2.25 / / / 82.8
CFG-only, $w=3$ 0.73 2.28 17.1 4.48 32.8 85.8
MoG-PG, $w_{1}=2.7,w_{2}=0.15$ 0.70 2.22 16.8 4.14 32.9 86.1
MoG-HG, $w_{1}=2.7,w_{2}=2.5,w_{3}=1.2$ 0.68 2.20 17.2 4.06 33.1 86.6
Table 4: Comparison between different guidance methods on T2A.
@@
Table 5: Comparison between different guidance methods on V2A.
@@
4.2 Main results
@@
T2A generation results. We conduct a comparison study of audio generation quality across GT (i.e., ground-truth audio) and a range of systems, including AudioGen, AudioGen-Large (Kreuk et al., 2022), Make-An-Audio (Huang et al., 2023b), TANGO-AF&AC-FT-AC (Kong et al., 2024), AudioLDM-Large-Full (Liu et al., 2023), AudioLDM 2, AudioLDM 2-Large (Liu et al., 2024a), and Stable Audio Open (Evans et al., 2025). The descriptions of these models are further detailed in the Appendix F.2. For AudioGen and AudioLDM, we report the metrics as presented in their original papers, and for the rest of the methods, we cite the results from ETTA (Lee et al., 2025). To demonstrate the effectiveness of our method, we further evaluate the base model using only CFG. The best results are achieved when the CFG scale is set to $w=7$. For a fair comparison, we fix the number of function evaluations (NFE) to 400  [^2] This corresponds to 200 sampling steps for CFG and 100 for HG. For PG, we instead keep the number of sampling steps consistent with HG, yielding an NFE of 300, as we observed that further increasing it leads to slight performance degradation for PG. for both our method and the CFG-only baseline, ensuring equal inference cost. All evaluations are conducted on the AudioCaps test set using standard objective metrics for quantitative comparison. The main results are summarized in Table 1. We have the following conclusions:
@@
Under equal inference resources, our method uniformly outperforms CFG-only. In terms of audio quality, our proposed methods demonstrate significantly improved performance compared to CFG-only under the same inference budget. Specifically, both PG and HG outperform CFG-only across all objective metrics. Our results show that improved quality can be achieved without increasing inference cost. Notably, our HG-improved variant gives an updated result, achieving a FAD of 1.38, KL of 1.44, IS of 13.58, and a CLAP score of 0.54.
@@
HG is better than PG. Furthermore, we observe that HG consistently outperforms PG. While both methods significantly surpass the CFG-only baseline, HG achieves better performance than PG across all evaluated metrics. For example, HG achieves a lower FAD and KL than PG, indicating better distributional fidelity. Moreover, the higher CLAP score indicates that HG achieves better text-audio alignment than PG. We hypothesize these improvements stem from the structured formulation of HG, which avoids the potential conflicts present in PG.
@@
V2A generation results. To further investigate the potential of AudioMoG, we also fine-tune the T2A model to V2A with CLIP features (Radford et al., 2021) and validate it on the VGGSound (Chen et al., 2020) test set. We evaluated our models using several metrics to assess audio quality, video-audio semantic alignment and temporal alignment, including FAD, KL, IS, FD, ImageBind Score (IBS) (Girdhar et al., 2023) and temporal alignment accuracy (Align acc) introduced in Diff-Foley (Luo et al., 2023). Then we compare our results with GT and a variety of strong systems, including IM2WAV (Sheffer & Adi, 2023), Diff-Foley (Luo et al., 2023), FoleyGen (Mei et al., 2024), VTA-LDM (Xu et al., 2024), FoleyCrafter (Zhang et al., 2024), V2A-Mapper (Wang et al., 2024a), VAB-Encodec (Su et al., 2024), and VATT (Liu et al., 2024b). The comparative results are summarized in Table 3. A comprehensive improvement across most metrics, including the critical issue of temporal alignment in V2A, demonstrates that our method uniformly enhances cross-modal audio generation. More details about the V2A experiment are provided in Appendix G.
@@
Text-to-music and image generation results. To further demonstrate the effectiveness of our proposed method, we also perform our approach on text-to-music generation with our DiT base model, and conditional image generation on the public EDM (Karras et al., 2024b) checkpoint. The results are shown in Appendix C and D respectively, which again verifies the efficacy of MoG across different tasks and modalities.
@@
Diverse samplers. In audio generation tasks, we use DPM++ solver, while in image generation, we use the Heun sampler, demonstrating the robustness of our methods on different samplers.
4.3 Additional results
@@
Figure 3: Performance comparison of HG and CFG-only across different NFEs. As shown, HG consistently outperforms CFG-only across all settings.
@@
Comparison across various NFEs. We compare the performance of HG and CFG-only under different numbers of function evaluations (NFE), each using their respective optimal settings. Specifically, we examine NFE values of 100, 200, 300 and 400, and report the results in Figure 3. Across all NFE levels, HG consistently outperforms CFG-only in terms of FAD and IS. This consistent superiority indicates that our method is more robust across varying computational budgets, maintaining high-quality generation even under stricter inference constraints. These results highlight the strong adaptability and scalability of HG, which delivers reliable quality improvements regardless of available inference resources, while also unlocking higher performance ceilings when additional computation is allowed.
@@
Comparison in the same guidance scale. To validate the effectiveness of our approach, we compare different guidance methods on both T2A and V2A. For NFE, we fix it to 300 for PG and 400 for the rest. For the guidance scale, we set $w_{1}=w_{3}=1$, $w_{3}=1$, $w_{1}=w_{2}=1$ and $w_{2}=1$ in the HG setting, which corresponds to no guidance, CFG-only, AG-only and PG, respectively. The remaining guidance scales are consistent with HG in Table 1 and Table 3 (AG-only also achieves the best performance at this scale). The results are shown in Table 5 and Table  5. We observe that HG and PG outperform no guidance, CFG-only and AG-only across FAD, KL and IS, indicating that combining guidance directions yields better performance than using either one alone. Furthermore, HG surpasses PG on both tasks, further confirming the advantage of the hierarchical guidance structure.
@@
Impact of guidance scales. We further investigate the influence of guidance scales in PG and HG on the audio generation quality. Our analysis reveals that these parameters have an effect on the balance between fidelity and diversity in the generated audio. Detailed experimental results and discussions are provided in Appendix B.
@@
Case study. Apart from the objective and subjective evaluations, we conduct a case study as shown in Figure 4. For each of the guidance methods, CFG, PG, and HG, we provide three test text prompts. As shown, CFG is prone to produce less structured results. In comparison, PG has shown improved quality and HG produces the strongest outcome. These results are consistent with our objective test results and the human evaluations. We provided more generation results in Appendix I.
Figure 4: Case study comparing the spectrogram outputs of different guidance strategies (CFG, PG, HG) under various text prompts. HG consistently demonstrates superior harmonic structure modeling and clearer spectral patterns compared to PG and CFG. While PG shows moderate improvements, CFG often struggles to capture harmonics and yields blurrier, less structured results, particularly for complex prompts. These examples visually highlight the effectiveness of hierarchical guidance in improving fidelity and temporal structure.
5 Related work
Diffusion-based cross-modal audio generation, such as T2A and V2A generation systems, primarily employ CFG during sampling, while it may suffer from sub-optimal synthesis quality as discussed. Recent work (Chidambaram et al., 2024) has theoretically justified the empirical finding that a large guidance scale which can ensure synthesis quality may result in reduced generation diversity. To address this issue, recent works (Karras et al., 2024a; Hong et al., 2023; Li et al., 2025; Sadat et al., 2024) carefully design different weaker models to replace the unconditional term in CFG, as noted in Jeon (2025). Among them, AG is representative as its general formulation, and has been extended to diverse data modalities and generation tasks (Phunyaphibarn et al., 2025; Jeon, 2025; Hyung et al., 2025; Lee et al., 2025). In this work, we propose a mixture-of-guidance framework, exploiting complementary advantages of different guidance methods, aiming at improving guided audio generation quality without sacrificing inference efficiency.
6 Conclusion

### v2

@@ IV AudioMoG
As discussed above, previous guidance methods, CFG and AG, guide the sampling process from distinct perspectives. CFG enhances consistency with conditional information, while AG removes dispersion by mitigating errors. The effectiveness of CFG has been extensively validated in audio generation across various data types [42, 14], conditional modalities such as text [18, 22] and video [46, 59], and network architectures [42, 14, 39, 24]. However, CFG can still yield suboptimal results due to its overemphasis on condition information. Recent work ETTA [38] explores AG on T2A generation but empirically observes limited synthesis quality and strong sensitivity to the choice of weak model. This suggests that the limitations of CFG remain unresolved in T2A tasks.
Motivated by these challenges, we propose a new guidance framework, making the first attempt to introduce the advantages of AG into audio generation. Our method, AudioMoG, introduces a mixture strategy to leverage the cumulative benefits of diverse guidance mechanisms, thus improving generation quality without sacrificing diversity. Specifically, AudioMoG allows CFG and AG to be combined either in parallel or hierarchically, forming two variants.
Parallel Guidance
As analyzed in Section III and shown in Fig. 2, CFG and AG result in complementary behaviors. To exploit these synergies, a straightforward mixture strategy is to employ them in parallel at each sampling step, which we call Parallel Guidance (PG for short). Specifically, the conditional predictor $\bm{\epsilon_{\theta}}(\bm{z}_{t},t,\bm{c})$ used in (3) is replaced by:
$\bm{\epsilon}_{\text{PG}}(\bm{z}_{t},t,\bm{c})=\bm{\epsilon_{\theta}}(\bm{z}_{t},t)+w_{1}(\bm{\epsilon_{\theta}}(\bm{z}_{t},t,\bm{c})-\bm{\epsilon_{\theta}}(\bm{z}_{t},t))$
$+w_{2}(\bm{\epsilon_{\theta}}(\bm{z}_{t},t,\bm{c})-\bm{\epsilon}_{\bm{\theta}_{\text{bad}}}(\bm{z}_{t},t,\bm{c})),$ (8)
where $\bm{\epsilon}_{\bm{\theta}_{\text{bad}}}(\bm{z}_{t},t,\bm{c})$ represents the weak version of conditional model (e.g., under-trained), serving as a reference to correct errors in $\bm{\epsilon_{\theta}}(\bm{z}_{t},t,\bm{c})$. This formulation unifies CFG and AG: setting $w_{2}=0$ recovers CFG (Equation (5)), and setting $w_{1}=0$ recovers AG (Equation (6)).
The first guidance term (scaled by $w_{1}$) reinforces alignment with the condition $\bm{c}$. However, excessive reliance on this term can lead to mode collapse, where generations become overly deterministic. In contrast, the second term (scaled by $w_{2}$) compares the strong and weak models under the same condition, steering the samples away from dispersed regions without skewing them further for condition compliance. Together, the two terms in PG provide complementary corrective signals.
Hierarchical Guidance
While PG offers a simple and effective integration of CFG and AG, it implicitly assumes their compatibility at each sampling step. However, as discussed in Section III, these two methods guide the model in potentially interfering directions: emphasizing the condition can lead to mode collapse, while correcting from weak references may introduce semantic drift. To better coordinate these behaviors, we present a more structured design, Hierarchical Guidance (HG for short), which applies guidance in a two-stage manner. Specifically, we first apply CFG to both strong and weak models, and then apply AG to their respective outputs [^1] HG in CFG-AG or AG-CFG orders yields equivalent guidance family, as proven in Appendix A:
@@
Here, a new cross-term $\bm{\epsilon}_{\bm{\theta}_{\text{bad}}}(\bm{z}_{t},t)$ (i.e., the unconditional bad version) is incorporated compared to PG. This hierarchical structure promotes a more principled interaction between condition- and model-aware signals. In contrast to PG, which only guides the good model using two distinct directions, HG also refines the bad model, yielding a more balanced corrective process. Moreover, HG generalizes PG. When $w_{2}=1$, (9) reduces exactly to (IV), demonstrating the broader expressiveness and adaptability of HG across different sampling regimes.
A potential concern with HG is increased computational cost per sampling step due to multiple model evaluations, as noted in ETTA [38]. Hence, we conduct a comprehensive comparison in Section V-C, matching inference time between CFG and HG. Results show that HG delivers overall better performance under the same inference budget.
V Experiments
V-A T2A experiment setup
Datasets
We mainly use AudioSet [17], FSD50k [15] and Clotho v2 [11]. To maintain consistency, each track in these databases was segmented into 10-second clips and resampled at 16 kHz. The details of these datasets are further introduced in the Appendix F-A. To compare with prior work, we evaluated our models on the widely used AudioCaps benchmark [33], which consists of about 1K 10-second audio clips.
Model configurations
We trained the base model for 1 M iterations with a batch size of 8 per GPU. We used the AdamW optimizer with a learning rate of 5e-5 and a condition drop $p_{\text{uncond}}=0.1$ for CFG. The bad model was trained under the same configuration as the main model, but with only 0.1 M iterations. During inference, we use DPM++ 2M SDE [45]. A more detailed description of the model configurations and compression networks is provided in the Appendix F-C.
Evaluation metrics
We conduct a comprehensive evaluation of our models using both objective and subjective evaluations to assess audio generation quality and text-audio alignment. Objective metrics include commonly used Fréchet Audio Distance (FAD), Kullback-Leibler (KL) divergence, Inception Score (IS), Fréchet Distance (FD), and LAION-CLAP score [57]. For subjective evaluation, we recruited 20 human raters to score two aspects: (i) overall perceptual quality (OVL), and (ii) semantic relevance to the input text (REL). Both scores are rated on a 1–5 scale. More details are introduced in Appendix F-E and F-F, respectively.
TABLE I: Objective metrics for T2A generation on AudioCaps test set. The best performance for each metric is highlighted in bold, while the second-best is marked with an underline.
@@ V-A T2A experiment setup
AudioGen [35] 3.13 2.09 / / /
AudioGen-Large [35] 1.82 1.69 / / /
Make-An-Audio [23] 1.61 1.61 7.29 18.32 /
TANGO-AF&AC-FT-AC [34] 2.54 / 11.04 17.19 /
AudioLDM-Large-Full [42] 1.96 1.59 8.13 23.31 0.43
AudioLDM 2 [43] 2.09 1.79 8.14 26.44 0.50
AudioLDM 2-Large [43] 1.89 1.54 8.55 26.18 0.53
Stable Audio Open [14] / 2.14 / / 0.35
CFG-only 1.76 1.44 13.46 20.94 0.54
MoG-PG (Ours) 1.54 1.47 13.47 18.50 0.53
MoG-HG (Ours) 1.38 1.44 13.58 18.87 0.54
TABLE II: Subjective metrics for T2A generation on AudioCaps samples.
@@
TABLE III: Objective metrics for V2A generation on VGGSound test set.
Model FAD $\downarrow$ KL $\downarrow$ IS $\uparrow$ FD $\downarrow$ IBS $\uparrow$ AA $\uparrow$
@@
IM2WAV [53] 6.41 2.54 / / 19.0 74.3
Diff-Foley [46] 5.79 3.12 10.8 21.90 20.4 89.9
FoleyGen [48] 1.65 2.35 / / 26.1 73.8
VTA-LDM [59] 2.01 2.37 10.4 12.80 26.2 77.0
FoleyCrafter [61] 2.32 2.54 9.9 18.10 27.7 83.6
V2A-Mapper [55] 0.90 2.68 12.5 8.35 22.4 78.3
VAB-Encodec [54] 2.69 2.58 / / / /
VATT w/o text [44] 2.35 2.25 / / / 82.8
CFG-only 0.73 2.28 17.1 4.48 32.8 85.8
MoG-PG (Ours) 0.70 2.22 16.8 4.14 32.9 86.1
MoG-HG (Ours) 0.68 2.20 17.2 4.06 33.1 86.6
TABLE IV: Comparison of different guidance methods.
(a) T2A results
@@
(b) V2A results
@@
V-B Main results
T2A generation results
@@ V-B Main results
We conduct a comparison study of audio generation quality across GT (i.e., ground-truth audio) and a range of systems. The descriptions of these models are further detailed in the Appendix F-B. For AudioGen and AudioLDM, we report the metrics as presented in their original papers, and for the rest of the methods, we cite the results from ETTA [38]. In MoG-PG we set $w_{1}=4.6,w_{2}=0.2$, while in MoG-HG, we set $w_{1}=4.0,w_{2}=3.3,w_{3}=1.2$. To demonstrate the effectiveness of our method, we further evaluate the base model using only CFG. The best results are achieved when the CFG scale is set to $w=7$. For a fair comparison, we fix the number of function evaluations (NFE) to 400222This corresponds to 200 sampling steps for CFG and 100 for HG. For PG, we instead keep the number of sampling steps consistent with HG, yielding an NFE of 300, as we observed that further increasing it leads to slight performance degradation for PG. for both our method and the CFG-only baseline, ensuring equal inference cost. All evaluations are conducted on the AudioCaps test set using standard objective metrics for quantitative comparison. The main results are summarized in Table I. We have the following conclusions:
@@
Under equal inference resources, our method uniformly outperforms CFG-only. In terms of audio quality, our proposed methods demonstrate significantly improved performance compared to CFG-only under the same inference budget. Specifically, both PG and HG outperform CFG-only across all objective metrics. Our results show that improved quality can be achieved without increasing inference cost.
@@
HG is better than PG. Furthermore, we observe that HG consistently outperforms PG. While both methods significantly surpass the CFG-only baseline, HG achieves better performance than PG across all evaluated metrics, indicating better distributional fidelity and text-audio alignment. We hypothesize these improvements stem from the structured formulation of HG, which avoids the potential conflicts present in PG.
V2A generation results
@@
To further investigate the potential of AudioMoG, we also fine-tune the T2A model to V2A with CLIP features and validate it on the VGGSound [4] test set. We evaluated our models using several metrics to assess audio quality, video-audio semantic alignment and temporal alignment, including FAD, KL, IS, FD, ImageBind Score (IBS) [19] and temporal alignment accuracy (AA) introduced in Diff-Foley [46]. Then we compare our results with GT and a variety of systems. The comparative results are summarized in Table III. A comprehensive improvement across all metrics, including the critical issue of temporal alignment in V2A, demonstrates that our method uniformly enhances cross-modal audio generation. More details about the V2A experiment are provided in Appendix G.
Text-to-music and image generation results
@@
To further demonstrate the effectiveness of our proposed method, we also perform our approach on text-to-music (T2M) generation with our DiT base model, and conditional image generation on the public EDM [30] checkpoint. The results are shown in Appendix C and D respectively, which again verifies the efficacy of MoG across different tasks and modalities.
Diverse samplers
@@
In audio generation tasks, we use DPM++ solver, while in image generation, we use the Heun sampler, demonstrating the robustness of our methods on different samplers.
V-C Additional results
@@
Figure 3: Performance comparison of HG and CFG-only under different NFEs.
Comparison across various NFEs
@@
We compare the performance of HG and CFG-only under different numbers of function evaluations (NFE), each using their respective optimal settings. Specifically, we examine NFE values of 100, 200, 300 and 400, and report the results in Fig. 3. Across all NFE levels, HG consistently outperforms CFG-only in terms of FAD and IS. This consistent superiority indicates that our method is more robust across varying computational budgets, maintaining high-quality generation even under stricter inference constraints. These results highlight the strong adaptability and scalability of HG, which delivers reliable quality improvements regardless of available inference resources, while also unlocking higher performance ceilings when additional computation is allowed.
Comparison in the same guidance scale
@@
To validate the effectiveness of our approach, we compare different guidance methods on both T2A and V2A. For NFE, we fix it to 300 for PG and 400 for the rest. For the guidance scale, we set $w_{1}=w_{3}=1$, $w_{3}=1$, $w_{1}=w_{2}=1$ and $w_{2}=1$ in the HG setting, which corresponds to no guidance, CFG-only, AG-only and PG, respectively. The remaining guidance scales are consistent with HG in Table I and Table III (AG-only also achieves the best performance at this scale). The results are shown in Table IV(a) and Table IV(b). The superiority of PG and HG indicates that combining guidance directions yields better performance than using either one alone. Furthermore, HG surpasses PG on both tasks, further confirming the advantage of the hierarchical structure.
Impact of guidance scales
@@
We further investigate the influence of guidance scales in PG and HG on the audio generation quality. Our analysis reveals that these parameters have an effect on the balance between fidelity and diversity in the generated audio. Detailed experimental results and discussions are provided in Appendix B.
Case study
@@
Apart from the objective and subjective evaluations, we conduct a case study for each of the guidance methods in Appendix I. As shown, CFG is prone to produce less structured results, while PG has shown improved quality and HG produces the strongest outcome. These results are consistent with our objective test results and the human evaluations.
VI Conclusion